# Pcc-tuning: Breaking the Contrastive Learning Ceiling in Semantic Textual Similarity

### v1

@@ 4.2 Main Results
In addition to challenges in fully harnessing fine-grained annotated data, another significant drawback of contrastive learning is the need for large batch sizes to prevent model collapse, which consumes substantial computational resources (Jiang et al., 2023b; Zhang et al., 2024). To explore the impact of batch size on Pcc-tuning’s performance, we conduct experiments detailed in Appendix A. The findings indicate that Pcc-tuning exhibits strong robustness to varying batch sizes. Additionally, we also assess Pcc-tuning on seven transfer tasks, with outcomes recorded in Appendix B.
@@ 4.3 Pcc-tuning under Various Prompts
In a pioneering effort to employ generative PLMs for embedding derivation, the Explicit One-word Limitation (EOL) format of the manual template, proposed by PromptEOL (Jiang et al., 2023b), has become the most widely adopted prompt in sentence representation research. Recently, Zhang et al. (2024) introduced two templates that deviate from the EOL structure, namely PromptSTH and PromptSUM. They demonstrated that adherence to the EOL format is not necessary for effective PLM fine-tuning. The specific forms of these prompts are depicted in Table 3, where [X] represents the input text, and the parts highlighted in red signify the positions from which the model extracts embeddings.
To further validate the versatility of our approach, we assess the average Spearman’s correlation scores across seven STS tasks using these prompts as the templates for both stages of Pcc-tuning. The corresponding results are delineated in Table 4. It can be seen that regardless of the chosen prompt, Pcc-tuning consistently enhances the model’s performance from approximately 85 to around 90, with minimal impact from the different templates on the final outcomes. This finding suggests when applying Pcc-tuning to downstream tasks, there is little need for laborious prompt searches, thereby offering significant application potential.
@@
| PLMs | Templates | Stage-1 | Stage-2 |
| OPT6.7b | PromptEOL | 85.52 | 89.29 |
| PromptSUM | 85.57 | 89.39 | |
| PromptSTH | 85.51 | 89.41 | |
| LLaMA7b | PromptEOL | 85.48 | 90.38 |
| PromptSUM | 85.47 | 90.13 | |
| PromptSTH | 85.40 | 90.29 | |
| LLaMA27b | PromptEOL | 85.40 | 90.32 |
| PromptSUM | 85.53 | 90.31 | |
| PromptSTH | 85.31 | 90.34 | |
| Mistral7b | PromptEOL | 85.50 | 90.39 |
| PromptSUM | 85.83 | 90.57 | |
| PromptSTH | 85.66 | 90.61 |
@@ 5 Related Work
Contrastive learning is currently the principal strategy employed by the NLP community for addressing STS tasks, and our method, Pcc-tuning, is specifically designed to overcome the inherent limitations of contrastive learning, particularly its inability to fully leverage the fine-grained annotated information in text pairs.
Prior to the rise of contrastive learning-based text representation schemes, Sentence-BERT had already proposed enhancing the semantic encoding capabilities of PLMs using the STS-B training set (Reimers and Gurevych, 2019). However, subsequent contrastive learning approaches such as SimCSE (Gao et al., 2021), PromptBERT (Jiang et al., 2022), and CoT-BERT (Zhang et al., 2023) have demonstrated superior performance across the seven STS benchmarks collected in SentEval, thereby making them the focal point of current academic research and development.
Among these efforts, RankCSE (Liu et al., 2023) also recognized that contrastive learning fails to capture the fine-grained ordinal relationships between texts and advocated for the use of Jensen-Shannon divergence to ensure rank consistency of embeddings derived under different dropout masks. However, this technique is only applicable in unsupervised scenarios. Supervised STS solutions, such as PromptEOL (Jiang et al., 2023b), still predominantly employ InfoNCE Loss to update model parameters, thus falling into the performance bottlenecks discussed in this paper.
To the best of our knowledge, this study is the first to propose and substantiate the performance upper bound of contrastive learning methods. Additionally, Pcc-tuning is the inaugural method capable of achieving Spearman’s correlation scores above 87 on standard STS tasks, marking a significant advancement in the field.
@@ 6 Conclusion
In this paper, we first analyze the structure of contrastive learning loss functions, highlighting that their coarse-grained categorization of semantic relationships between text pairs renders contrastive learning akin to a binary classifier. Building on this insight, we rigorously derive the optimal Spearman correlation achievable by a binary classifier in STS tasks, demonstrating that the upper bound for the Spearman’s correlation score of contrastive learning methods is 87.5. This finding effectively explains the performance bottlenecks encountered by current sentence representation methods in STS tasks.
To achieve further breakthroughs, we introduce Pcc-tuning, a strategy that effectively harnesses fine-grained annotated information. Pcc-tuning leverages a two-stage training pipeline and utilizes Pearson’s correlation coefficient as the loss function in the second stage to fully exploit the ordinal relationships between text pairs. Extensive experimental results demonstrate that Pcc-tuning significantly enhances the quality of the generated embeddings, and this improvement is consistently observed across different PLMs, prompts, and batch sizes.
@@
In preparing the training dataset for the second stage of Pcc-tuning, we employ a mixed corpus composed of the training sets from STS-B and SICK-R. However, the label scales of these two datasets are not completely congruent. Specifically, the STS-B training set contains 5,749 text pairs with similarity scores spanning from 0 to 5, whereas the SICK-R training set includes 4,500 text pairs with similarity scores ranging from 1 to 5. To unify their annotation scales, we transform each label in the SICK-R training set using the formula , thereby converting the labels to the range. Given that this transformation is a simple linear mapping, it is likely that some vital manually annotated information is lost, potentially hindering Pcc-tuning from reaching its optimal performance on the evaluation benchmarks.
Appendix A Pcc-tuning under Different Batch Sizes
Here, we explore the impact of batch size on the performance of Pcc-tuning. In line with previous sections, we employ four 7B-scale generative PLMs as backbones and report the average Spearman’s correlation scores of Pcc-tuning across seven STS tasks in SentEval under various parameter configurations. We continue to utilize PromptSTH as the manual template for encapsulating input sentences, which is also the default setting for Pcc-tuning.
@@ Appendix A Pcc-tuning under Different Batch Sizes
Table 5 presents the results from these experiments. Despite the significant differences between batch sizes of 192 and 256, the resulting Spearman’s correlation scores are remarkably similar, with both maintaining high performance levels. This observation indicates that Pcc-tuning is not sensitive to batch size. Further combined with the findings from Section 4.3, where Pcc-tuning exhibits minimal performance fluctuations under different prompts, it can be concluded that Pcc-tuning possesses exceptional robustness and can easily adapt to a variety of hyperparameter configurations.
@@
| OPT6.7b | 192 | 89.34 |
| 216 | 89.41 | |
| 224 | 89.38 | |
| 256 | 89.35 | |
| LLaMA7b | 192 | 90.26 |
| 224 | 90.28 | |
| 232 | 90.29 | |
| 256 | 90.25 | |
| LLaMA27b | 192 | 90.34 |
| 200 | 90.32 | |
| 216 | 90.30 | |
| 256 | 90.32 | |
| Mistral7b | 192 | 90.54 |
| 224 | 90.61 | |
| 232 | 90.60 | |
| 256 | 90.54 |
| Model | MR | CR | SUBJ | MPQA | SST2 | TREC | MRPC | Avg. |
| GloVe | 77.25 | 78.30 | 91.17 | 87.85 | 80.18 | 83.00 | 72.87 | 81.52 |
| Skip-thought | 76.50 | 80.10 | 93.60 | 87.10 | 82.00 | 92.20 | 73.00 | 83.50 |
| Avg. BERT | 78.66 | 86.25 | 94.37 | 88.66 | 84.40 | 92.80 | 69.45 | 84.94 |
| BERT-CLS | 78.68 | 84.85 | 94.21 | 88.23 | 84.13 | 91.40 | 71.13 | 84.66 |
| IS-BERT | 81.09 | 87.18 | 94.96 | 88.75 | 85.96 | 88.64 | 74.24 | 85.83 |
| SimCSE-BERT | 81.18 | 86.46 | 94.45 | 88.88 | 85.50 | 89.80 | 74.43 | 85.81 |
| PromptBERT | 80.74 | 85.49 | 93.65 | 89.32 | 84.95 | 88.20 | 76.06 | 85.49 |
| Implementation on | ||||||||
| Pcc-tuning | 89.40 | 92.77 | 95.95 | 91.29 | 94.34 | 95.80 | 76.00 | 90.79 |
| Implementation on | ||||||||
| Pcc-tuning | 89.59 | 92.74 | 95.63 | 90.19 | 94.12 | 95.00 | 77.62 | 90.70 |
| Implementation on | ||||||||
| Pcc-tuning | 89.72 | 93.51 | 95.95 | 90.75 | 94.34 | 94.60 | 76.12 | 90.71 |
| Implementation on | ||||||||
| Pcc-tuning | 88.78 | 92.21 | 95.77 | 89.45 | 93.74 | 95.80 | 74.09 | 89.98 |
Appendix B Transfer Tasks
In addition to the standard STS benchmarks, we also evaluate Pcc-tuning on several transfer tasks, including MR, CR, SUBJ, MPQA, SST2, TREC, and MRPC. The results, displayed in Table 6, demonstrate that Pcc-tuning consistently outperforms the baselines across all datasets. Notably, its average score exceeds those of SimCSE and PromptBERT by 4 to 5 percentage points, underscoring the ability of Pcc-tuning to generate high-quality embeddings applicable across a broad range of scenarios.

### v2

@@ 4.2 Main Results
In addition to the inability in fully harnessing fine-grained annotated data, another significant drawback of contrastive learning is its reliance on large batch sizes to ensure negative sample diversity, which consumes substantial computational resources (Jiang et al., 2023b; Zhang et al., 2024b). To explore Pcc-tuning’s memory consumption and the impact of batch size on model performance, we conducted relevant experiments detailed in Appendix D. The findings indicate that Pcc-tuning demonstrates superior memory efficiency and exhibits strong robustness to varying batch sizes.
4.3 Pcc-tuning vs. Two-Stage Contrastive Learning
This section addresses an intriguing question: Can contrastive learning, when supplemented with fine-grained annotated data, further enhance model performance? Specifically, when employing the filtered STS-B and SICK-R training sets for two-stage parameter updates, does Pcc-tuning outperform continuous contrastive learning?
As analyzed in section 2, contrastive learning methods are limited in their ability to fully leverage fine-grained annotated data. Therefore, to apply the STS-B and SICK-R training sets to contrastive learning models, a threshold must be selected to identify suitable positive sample pairs, which inevitably leads to significant data loss. After balancing dataset scale and quality, we chose to treat text pairs with similarity scores greater than 4.0 as positive samples. Following this step, the original 5,398 data pairs were reduced to 1,543.
Table 3 presents the results of the above experiments. In the second column, "Contrastive" and "Pearson" refer to fine-tuning the first-stage checkpoint using either contrastive learning or Pcc-tuning, respectively. The "Performance" column reports the model’s average Spearman’s correlation scores across seven STS benchmarks. The results clearly show that continuing with contrastive learning not only yields significantly inferior results compared to Pcc-tuning, but also underperforms the model’s first-stage outcomes.
| PLMs | Strategy | Performance |
| OPT6.7b | Stage I | 85.57 |
| Contrastive | 77.29 | |
| Pearson | 86.93 | |
| LLaMA7b | Stage I | 85.47 |
| Contrastive | 79.47 | |
| Pearson | 87.35 | |
| LLaMA27b | Stage I | 85.53 |
| Contrastive | 85.38 | |
| Pearson | 87.80 | |
| Mistral7b | Stage I | 85.83 |
| Contrastive | 75.47 | |
| Pearson | 87.86 |
These findings are not surprising, as contrastive learning methods require large batch sizes to avoid model collapse. Indeed, most mainstream contrastive learning-based text representation models employ batch sizes of 256 or more. In comparison, such a small amount of annotated data is hardly sufficient to effectively support contrastive learning. However, it is important to note that due to the coarse-grained semantic partitioning inherent in InfoNCE Loss, even with a larger corpus, contrastive learning methods cannot surpass Pcc-tuning. This is because Pcc-tuning possess higher data utilization efficiency and is more closely aligned with the evaluation metrics of STS tasks.
@@ 4.4 Pcc-tuning under Various Prompts
The Explicit One-word Limitation (EOL) template, introduced by PromptEOL (Jiang et al., 2023b), represents a pioneering effort in employing generative PLMs for embedding derivation and has become the most widely adopted prompt in sentence representation research. Recently, Zhang et al. (2024b) proposed two alternative templates, PromptSTH and PromptSUM, which deviate from the EOL structure. Their findings demonstrated that strict adherence to the EOL format is not necessary for effective PLM fine-tuning. The specific forms of these prompts are depicted in Table 4, where [X] represents the input text, and the parts highlighted in red denote the positions from which the model extracts embeddings.
@@
To further validate the versatility of our approach, we assessed the average Spearman’s correlation scores on seven STS tasks using these prompts as the templates for both stages of Pcc-tuning. The corresponding results are delineated in Table 5. As evidenced by the results, Pcc-tuning consistently improves model performance, with minimal impact from the different templates on the final outcomes. This suggests that when applying Pcc-tuning to downstream tasks, there is little need for laborious prompt searches, thereby offering significant practical benefits.
| PLMs | Templates | Stage I | Stage II |
| OPT6.7b | PromptSTH | 85.51 | 86.86 |
| PromptEOL | 85.52 | 86.96 | |
| PromptSUM | 85.57 | 86.93 | |
| LLaMA7b | PromptSTH | 85.40 | 87.54 |
| PromptEOL | 85.48 | 87.67 | |
| PromptSUM | 85.47 | 87.35 | |
| LLaMA27b | PromptSTH | 85.31 | 87.64 |
| PromptEOL | 85.40 | 87.75 | |
| PromptSUM | 85.53 | 87.80 | |
| Mistral7b | PromptSTH | 85.66 | 87.70 |
| PromptEOL | 85.50 | 87.72 | |
| PromptSUM | 85.83 | 87.86 |
@@ 5 Related Work
Contrastive learning is currently the principal strategy within the NLP community for addressing STS tasks, and our proposed method, Pcc-tuning, is specifically designed to overcome the inherent limitations of these approaches.
Prior to the rise of contrastive learning-based text representation schemes, Sentence-BERT had already introduced the idea of enhancing the semantic encoding capabilities of PLMs using the STS-B training set (Reimers and Gurevych, 2019). However, subsequent contrastive learning methods, such as SimCSE (Gao et al., 2021), PromptBERT (Jiang et al., 2022), and CoT-BERT (Zhang et al., 2024a) have demonstrated superior performance across the seven STS benchmarks collected in SentEval, thereby making them the focal point of recent academic research and development.
To the best of our knowledge, this study is the first to propose and substantiate the theoretical performance upper bound of contrastive learning methods. Additionally, Pcc-tuning is the inaugural method capable of achieving Spearman’s correlation scores above 87 on standard STS tasks, marking a significant advancement in the field.
@@ 6 Conclusion
In this paper, we first analyze the structure of contrastive learning loss functions, highlighting that their coarse-grained categorization of semantic relationships among texts renders contrastive learning akin to a binary classifier. Building on this insight, we rigorously derive the optimal Spearman correlation achievable by a binary classifier in STS tasks, establishing that the upper bound for the Spearman correlation of contrastive learning methods is 87.5. To achieve further breakthroughs, we introduce Pcc-tuning, a novel strategy that effectively harnesses fine-grained annotated information. Pcc-tuning leverages a two-stage training pipeline and utilizes Pearson’s correlation coefficient as the loss function to fully exploit the ordinal relationships between text pairs. Extensive experimental results demonstrate that Pcc-tuning significantly enhances the quality of generated embeddings, with consistent performance gains observed across various PLMs, prompts, and batch sizes.
@@
In preparing the training dataset for the second stage of Pcc-tuning, we employ a mixed corpus composed of the training sets from STS-B and SICK-R, while filtering out any overlapping samples with the test sets. However, the labeling scales of these two datasets are not congruent. Specifically, the similarity scores in the STS-B training set span from 0 to 5, whereas the scores in the SICK-R training set range from 1 to 5. To unify their annotation scales, we transform each label in the SICK-R training set using the formula , thereby converting the scores to the range of . Given that this is merely a simple linear mapping, it is likely that some vital manually annotated information is lost, potentially hindering Pcc-tuning from reaching its optimal performance on the evaluation benchmarks.
Appendix A Derivation Details
Appendix B Data Filtering Method
In section 3, we mentioned that the fine-tuning data in the second stage of Pcc-tuning is derived from filtered STS-B and SICK-R training sets, aimed at preventing the model from encountering text pairs present in the evaluation benchmarks during parameter updates. Here, we provide a detailed description of the filtering process.
In standard STS tasks, each sample consists of two text strings, and , along with a floating-point number gs indicating the semantic similarity score between them. In our experimental setup, for any sample from the STS-B or SICK-R training sets, if a text pair exists in the test sets of STS12-16, STS-B, or SICK-R, where and , or and , we treat them as duplicates, regardless of whether and are identical. All duplicate training samples are then removed from the model’s fine-tuning corpus.
This process involves a highly stringent filtering mechanism. Under this approach, even the train and test sets of SICK-R itself contain overlapping samples (despite differences in their gs values). The original STS-B and SICK-R training sets consist of 5,749 and 4,500 samples, respectively. After filtering, they are reduced to 991 and 4,407 samples, respectively, resulting in a total of 5,398 text pairs. Moreover, as noted in the Limitations section of this paper, the annotation scales of these two datasets are not consistent, which leads to some information loss during the merging process. The scarcity of annotated data, combined with differences in scoring standards, posed additional challenges for this study.
Despite these challenges, Pcc-tuning still demonstrated strong performance (section 4, Appendix D). This suggests that in downstream scenarios with more abundant task-specific data, the advantages of Pcc-tuning over contrastive learning could become even more pronounced, highlighting its broad potential for application.
Appendix C Transfer Tasks
In the previous sections, we have thoroughly validated the exceptional performance of Pcc-tuning in semantic matching across seven well-established STS benchmarks. To further clarify its generalization capabilities, this section evaluates the transferability of Pcc-tuning through zero-shot testing on a variety of downstream tasks.
| LLaMA2 (raw) | PromptSUM | Pcc-tuning | |
| Banking77Classification | 56.38 | 85.41 | 86.09 |
| TwitterSemEval2015 | 79.05 | 87.66 | 87.73 |
| AskUbuntuDupQuestions | 41.46 | 58.00 | 61.19 |
| StackOverflowDupQuestions | 24.44 | 44.93 | 47.65 |
| CQADupstackEnglishRetrieval | 0.17 | 32.31 | 34.19 |
| LegalSummarization | 7.49 | 66.20 | 68.31 |
| FaithDial | 0.52 | 25.54 | 28.28 |
| PIQA | 1.06 | 30.81 | 31.07 |
We selected eight tasks of different types from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022) to cover four major areas: Retrieval, Reranking, Classification, and Pair Classification. For these experiments, we employed LLaMA27b as the backbone and directly loaded model checkpoints without any additional parameter updates. In other words, the PromptSUM and Pcc-tuning models used here are identical to those in Table 2.
Table 6 summarizes the results of these tests. In this table, "LLaMA2 (raw)" refers to the original LLaMA2 model without any integrated prompts. It can be observed that the transferability of raw LLaMA2 output vectors is quite poor, with scores on some tasks falling below 1%. This suggests a substantial gap between auto-regressive language modeling and effective embedding generation. However, the scores for PromptSUM demonstrate a significant improvement over LLaMA2 (raw), indicating that contrastive learning can enhance the embedding quality of generative PLMs. Pcc-tuning further amplifies this effect. Despite introducing only an additional 5,398 fine-grained annotated text pairs, Pcc-tuning consistently outperforms PromptSUM across multiple STS benchmarks and downstream tasks. These results highlight the strong generalizability of our proposed method and its effectiveness in various application scenarios.
Appendix D Memory Usage and Batch Sizes
Here, we examine the memory consumption of Pcc-tuning and analyze the impact of batch size on model performance. For implementation, we utilized four 24GB NVIDIA GPUs, with PromptSUM as the manual template across all experimental groups.
Table 7 presents the memory usage for each of the two fine-tuning stages of Pcc-tuning across different backbone models. The results show that optimizing the model with Pearson’s correlation coefficient in the second stage requires significantly fewer computational resources compared to contrastive learning in the first stage. Furthermore, given that the maximum sequence length supported in the second stage of Pcc-tuning is twice that of the contrastive learning stage, our method demonstrates a clear advantage in memory efficiency.
| PLMs | Stage | Memory (GB) |
| OPT6.7b | I | 91.83 |
| II | 58.44 | |
| LLaMA7b | I | 93.82 |
| II | 65.22 | |
| LLaMA27b | I | 93.82 |
| II | 69.37 | |
| Mistral7b | I | 93.41 |
| II | 73.31 |
Several factors contribute to this improvement, with batch size being a critical one. In standard InfoNCE Loss, negative instances for the current sample are drawn from other texts within the same batch. As a result, contrastive learning-based STS solutions typically require large batch sizes to provide sufficient reference information for optimizing embeddings. For example, SimCSE employs a batch size of 512 in supervised settings. Additionally, our experiments in section 4.3 also support this observation.
Given this context, we were interested in exploring how varying batch sizes affect performance when using Pearson’s correlation coefficient for training. Therefore, we tested Pcc-tuning on seven STS tasks collected in SentEval, using four 7B-level generative PLMs under different batch size conditions. Due to the relatively slow inference speed of 7B-scale models, we did not perform an exhaustive grid search, but intuitively selected several batch sizes for testing, which was sufficient to illustrate the key trends.
@@ Appendix D Memory Usage and Batch Sizes
The results are summarized in Table 8. It is evident that even with variations in batch size by several dozen, Pcc-tuning maintains consistently high performance. This indicates that our proposed method is not sensitive to batch size. Combined with the findings from section 4.4, where Pcc-tuning exhibits minimal performance fluctuations under different prompts, we conclude that Pcc-tuning demonstrates exceptional robustness and can easily adapt to a wide range of hyperparameter configurations.
@@
| OPT6.7b | 192 | 86.88 |
| 200 | 86.93 | |
| 240 | 86.90 | |
| 248 | 86.89 | |
| LLaMA7b | 192 | 87.34 |
| 200 | 87.35 | |
| 216 | 87.27 | |
| 224 | 87.29 | |
| LLaMA27b | 200 | 87.77 |
| 208 | 87.73 | |
| 216 | 87.80 | |
| 240 | 87.68 | |
| Mistral7b | 200 | 87.76 |
| 208 | 87.86 | |
| 216 | 87.75 | |
| 256 | 87.73 |